### v2

# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

Abstract
With the proliferation of spatio-textual data, Top-k KNN spatial keyword queries (TkQs), which return a list of objects based on a ranking function that considers both spatial and textual relevance, have found many real-life applications. To efficiently handle TkQs, many indexes have been developed, but the effectiveness of TkQ is limited. To improve effectiveness, several deep learning models have recently been proposed, but they suffer severe efficiency issues and there are no efficient indexes specifically designed to accelerate the top-k search process for these deep learning models.
To tackle these issues, we propose a novel technique, which Learns to Index the Spatio-Textual data for answering embedding based spatial keyword queries (called LIST). LIST is featured with two novel components. Firstly, we propose a lightweight and effective relevance model that is capable of learning both spatial and textual relevance. Secondly, we introduce a novel machine learning based Approximate Nearest Neighbor Search (ANNS) index to accelerate the top-k search process of the proposed relevance model, which utilizes a new learning-to-cluster technique to group relevant queries and objects together while separating irrelevant queries and objects. There exist two key challenges in building an effective and efficient index, i.e., the absence of high-quality labels and the unbalanced clustering results. We develop a novel pseudo-label generation technique to address the two challenges. Experimental results show that LIST significantly outperforms state-of-the-art methods on effectiveness, with improvements up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10, and is three orders of magnitude faster than the most effective baseline.
1 Introduction
With the proliferation of mobile Internet, spatio-textual (a.k.a geo-textual) data is being increasingly generated. Examples of spatio-textual data include (1) web pages with geographical information; (2) user-generated text content with location information, such as geo-tagged tweets and reviews related to local stores; (3) Points of Interest (POI) in local business websites or location-based apps DBLP:journals/pvldb/ChenLCLBLGZ21 ; (4) multimedia data containing both text and geo-locations, like photos shared on social platforms that provide both textual descriptions and geographic location. Meanwhile, with the prevalence of smartphones, accessing and querying spatio-textual data has become increasingly frequent. This trend calls for techniques to process spatial keyword queries efficiently and effectively, which take query keywords and location as input and return objects that match the given requirements. An example query is to search for a ‘delicious Italian restaurant’ that is close to the user’s hotel.
Spatial keyword queries have applications in many real-world scenarios such as geographic search engines, location-based services, and local web advertising tailored to specific regions. To meet diverse user needs, various spatial keyword queries have been introduced congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:journals/vldb/ChenCCJ21 . Among them, the Top-k KNN Spatial Keyword Query (TkQcongEfficientRetrievalTopk2009 retrieves top-k geo-textual objects according to a ranking function that considers both textual and spatial relevance. Specifically, TkQ computes textual relevance by traditional information retrieval models such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and uses a linear function of distance between query location and object location to evaluate spatial relevance (formulated in Section 3.1). According to the experimental evaluation liu2023effectiveness , TkQs return more relevant objects compared to several other spatial keyword queries like the Boolean KNN Query cary2010efficient . Most of the existing studies congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 on spatial keyword queries focus on improving the efficiency of handling spatial keyword queries. As such, various indexes DBLP:journals/vldb/ChenCCJ21 and corresponding query processing algorithms have been developed.
Although many indexes DBLP:journals/vldb/ChenCCJ21 have been developed to expedite the top-k search process of TkQs, the effectiveness of TkQs is limited. As discussed earlier, TkQ typically uses traditional models such as BM25 to compute textual relevance, but these models rely on exact word matching to evaluate textual relevance and thus suffer from the word mismatch issue, e.g., synonyms that consist of different tokens may convey the same meanings, which limits their effectiveness (detailed in Section 3.2). For example, when a user searches for ‘nearby bar’, nearby bars named ‘pub’ or ‘tavern’ will be ignored by these traditional models. Another limitation arises from using the linear function of distance to calculate spatial relevance. This practice is based on an assumption that the user’s spatial preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and does not hold on our real-life datasets, where the spatial preference for geo-textual objects exhibits a significant (non-linear) decrease with the increase of distance. This mismatch with real-world situations results in limited effectiveness (as to be detailed in Section 3.2).
To improve the effectiveness of TkQ, several deep learning based relevance models liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 have recently been proposed, but they suffer from severe efficiency issue. To the best of our knowledge, there are no efficient indexes specifically designed to expedite the top-k search process for these deep relevance models. For example, the recently proposed DrW liu2023effectiveness employs the BERT model devlin2018bert to generate word embeddings, identifies top-k relevant terms from object description for each query keyword based on the word embeddings, and utilizes an attention mechanism to aggregate the relevance scores of every query keyword as the textual relevance. It also uses the linear function of distance as the spatial relevance. The interactions between word embeddings bring very high query latency. And it relies on interactions between the word embeddings of queries and objects, which cannot be pre-computed until the query-object pairs are given, therefore there is no efficient index designed for it (detailed in Section 3.2).
To address the aforementioned limitations, this work aims to (1) design a lightweight yet effective relevance model, and (2) design an index to expedite the top-k search process of the proposed relevance model.
For the first objective, we propose a lightweight yet effective relevance model for embedding based spatial keyword queries, which capture the semantic meaning of query keywords in one high-dimensional embedding. The relevance model comprises three modules: a textual dual-encoder module, a novel learning based spatial relevance module, and a new weight learning module. The dual-encoder module converts query keywords and object descriptions into two separate embeddings and evaluates textual relevance based on the two embeddings. To overcome the limitation of linear distance function, as explained earlier, we propose a new learning-based spatial relevance module, which is capable of learning from real-world datasets to better evaluate spatial relevance. Additionally, a weight learning module is proposed to adaptively learn a weight to balance the textual and spatial relevance.
For the second objective, existing geo-textual indexes are developed based on traditional models, and cannot be used to handle embedding based spatial keyword queries. Although Approximate Nearest Neighbor Search (ANNS) indexes qin2020similarity ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 are designed for embedding retrieval, these ANNS indexes do not consider the spatial factor, which is essential for spatial keyword queries. Directly using these ANNS indexes for embedding based spatial keyword queries results in severe degradation of effectiveness (as to be shown in Section 5.2). An intuitive idea is to adapt these ANNS indexes to incorporate the spatial factor. For instance, the IVF index jegouProductQuantizationNearest2011 clusters embeddings using the K-means algorithm and route queries to a small number of close clusters to reduce the search space. However, incorporating spatial factors into the K-means algorithm is challenging as it is hard to set the weight to balance spatial and embedding similarities during index construction. Manually setting the weights is not only laborious but also ineffective, resulting in inferior effectiveness (as to be shown in Section 5.2). It is still an open problem to design an effective and efficient ANNS index to support embedding based spatial keyword queries.
To fill this gap, we develop a novel machine learning based ANNS index. To cluster both spatial and textual relevant objects together without manually setting the weight between spatial relevance and textual relevance, our proposed index utilizes the learning-to-cluster technique, which was originally developed for image clustering, to learn from pairwise relevant and irrelevant query-object pairs, thereby clustering relevant objects and queries together while separating the irrelevant ones. Existing learning-to-cluster studies Hsu16_KCL ; Hsu18_L2C demonstrate that high-quality pairwise similar/dis-similar labels are essential for training. Although these pairwise labels are easy to obtain for images, such pairwise positive labels between queries and objects in our problem are very sparse , and high-quality negative labels are absent. Additionally, when the number of clusters is large, the existing learning-to-cluster technique will produce a highly skewed cluster distribution Hsu16_KCL , which will hurt the index’s efficiency if the index is built on the cluster results. To address these challenges, we propose a novel pseudo-negative label generation method, which employs the trained relevance model to produce high-quality pseudo-negative labels. Through these informative labels, our index learns to precisely cluster relevant queries and objects together while separating the irrelevant ones, thereby producing a precise and balanced clustering result and constructing an effective and efficient ANNS index.
Based on the two proposed techniques, namely the lightweight yet effective relevance model and the machine learning based ANNS index, we propose a new solution, which learns to index spatio-textual data for answering embedding based spatial keyword queries (called LIST). LIST employs the proposed relevance model to pre-compute embeddings for geo-textual objects and uses the proposed index to cluster the relevant objects together while separating irrelevant ones. Given embedding based spatial keyword queries, LIST routes queries to a subset of clusters to reduce the search space. Subsequently, the relevance model is applied to re-rank the retrieved objects within the relevant clusters (detailed in Section 4).
The contributions of this work are summarized as follows:
  • New Relevance Model. We develop a lightweight yet effective relevance model. It is featured with a dual-encoder module to convert query keywords and object descriptions into two separate embeddings and compute textual relevance based on the two embeddings, a novel learning-based spatial relevance module to learn from real-word datasets to better evaluate the spatial relevance, and a new weight learning mechanism to determine the weights between textual and spatial relevance.
  • New Index. We develop a novel machine learning based ANNS index that is tailored for embedding based spatial keyword queries and accelerating the top-k search process of the proposed deep relevance model. Different from existing ANNS indexes, it utilizes the learning-to-cluster technique to cluster relevant objects and queries together while separating the irrelevant ones. To build an effective and efficient index, we propose a novel pseudo-label generation approach. To the best of our knowledge, this is the first index designed for embedding based spatial keyword queries and deep relevance model. Additionally, this is the first geo-textual index that employs neural networks for retrieval without relying on an explicit tree structure.
  • Extensive Experiments. We extensively evaluate the effectiveness and efficiency of our proposed LIST solution on three real-world datasets. Experimental results show that LIST significantly outperforms the state-of-the-art baseline algorithms for effectiveness by an improvement up to 19.21% and 12.79% in terms of NDCG@1 and Recall@10 respectively and is three orders of magnitude faster than the most effective baseline.
2 Related Work
Spatio-Textual Relevance Models. Spatial keyword queries have attracted extensive attention and many types of spatial keyword queries congEfficientRetrievalTopk2009 ; DBLP:conf/sigmod/ChenSM06 ; cary2010efficient are proposed. Among them, Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 aims to retrieve top-k geo-textual objects based on a ranking function that evaluates both spatial and textual relevance. Specifically, TkQ computes the textual relevance with traditional retrieval models like BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 and exploits a simple linear function of distance for spatial relevance. However, these traditional methods have limited effectiveness.
To enhance the effectiveness, several deep learning based methods DBLP:conf/aaai/ZhaoPWCYZMCYQ19 ; liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 have been developed for query-POI matching, which is essentially spatial keyword query. PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 considers geographic information by using location embedding techniques and combines it with textual word semantic representations for query-POI matching. DrW liu2023effectiveness computes the deep textual relevance on the term level of query keywords and object descriptions. It uses the attention mechanism to aggregate the scores of each term for spatial keyword queries and design a learning-based method to learn a query-dependent weight to balance textual and spatial relevance. MGeo DBLP:conf/sigir/DingCXHLZX23 employs a geographic encoder and a multi-modal interaction module, treating geographic context as a new modality and using text information as another modality. MGeo aligns these two modalities into the same latent space and computes relevance scores based on the produced representations. However, these deep learning based methods focus on improving the effectiveness but ignore the efficiency issue.
Different from previous spatio-textual relevance models that rely on word embeddings and complex models to compute textual relevance, our proposed model employs a dual encoder module to transform query keywords and object descriptions into two embeddings, computes textual relevance based on two embeddings, and we also propose a new learning-based spatial module to learn from real-world datasets to better estimate spatial relevance.
Spatio-Textual Indexes. Various spatio-textual indexes DBLP:conf/sigmod/ChenSM06 ; cary2010efficient ; DBLP:conf/icde/FelipeHR08 ; congEfficientRetrievalTopk2009 ; DBLP:journals/tkde/LiLZLLW11 ; DBLP:conf/ssd/RochaGJN11 ; DBLP:conf/icde/ZhangZZL13 ; tao2013fast ; vaid2005spatio ; DBLP:conf/cikm/GobelHNB09 ; DBLP:conf/cikm/ZhouXWGM05 ; DBLP:conf/sigmod/LuLC11 ; DBLP:conf/edbt/ZhangTT13 ; DBLP:conf/icde/ZhangOT10 ; DBLP:conf/icde/ZhangCMTK09 ; DBLP:journals/pacmmod/Sheng0F00C023 ; DBLP:conf/gis/Liu022 have been designed to efficiently answer spatial keyword queries. However, they are all designed for traditional retrieval models (e.g., TF-IDF), and are unsuitable for accelerating the top-k search process of these deep learning based methods.
[…]
Additionally, there exists a learned geo-textual index called WISK DBLP:journals/pacmmod/Sheng0F00C023 , which utilizes query workloads to partition geo-textual data to build a tree-based index and employ reinforcement learning techniques to optimize the index. However, WISK is designed for spatial keyword range queries, which treats query keywords and query region as Boolean filters, only retrieves objects containing all the query keywords within the given query region, and is not designed to support other queries like the TkQ, which uses keywords to compute textual relevance.
Different from previous spatio-textual indexes developed to expedite the top-k search for a ranking function that uses traditional relevance models like BM25 and TF-IDF, which depend on exact word matching to compute textual relevance, our goal is to design an index that accelerates the top-k search for a deep learning based relevance model, which relies on embeddings to compute textual relevance. It is still an open problem for geo-textual data.
Deep Textual Relevance Models and Approximate Nearest Neighbor Search Indexes. Our work is related to deep textual relevance models and the corresponding Approximate Nearest Neighbor Search (ANNS) index techniques. The deep textual relevance models can be broadly classified into two categories: interaction-focused models and representation-focused models guo2020deep . The first category of models (e.g., ARC-II DBLP:conf/nips/HuLLC14 and MatchPyramid DBLP:conf/aaai/PangLGXWC16 ) calculates the word-level similarity between queries and documents for textual relevance. DrW liu2023effectiveness belongs to this category. Although this category of methods may have better effectiveness, these methods are usually computationally expensive. The second category of models (e.g., DSSM huang2013learning ) extracts global semantic representation for input text and uses functions like the inner product to compute the relevance score between representations. This category of models used to be less effective than the first category of models. However, with the emergence of the dual-encoder architecture DBLP:conf/emnlp/KarpukhinOMLWEC20 ; DBLP:conf/icml/GuuLTPC20 , which employs Pre-trained Large Language Models (PLMs) to extract global semantic representation, it has become a well-established paradigm in document retrieval DBLP:journals/tois/GuoCFSZC22 ; DBLP:journals/corr/abs-2211-14876 ; ECML_23_zhou2023master_dense_retrieval ; SIGIR_23_wen2023offline_dense_retrieval ; DBLP:conf/naacl/QuDLLRZDWW21 . For the second category of methods, to support efficient online retrieval, the learned sentence embeddings are usually pre-computed offline and indexed by the ANNS indexes guo2020deep . Note that, models in the first category like DrW liu2023effectiveness are not efficient for online computation and retrieval since the complex word interaction function cannot be pre-calculated until we see the input query-object pairs according to guo2020deep .
ANNS indexes are developed to expedite the top-k search on high-dimensional embeddings. These techniques can be broadly categorized into two types. The first type focuses on searching a subset of the database. The representative methods include inverted file index (IVF) based methods DBLP:journals/pami/BabenkoL15 ; jegouProductQuantizationNearest2011 ; wang2020deltapq , hashing-based methods luoSurveyDeepHashing2023 ; liu2014sk ; liu2019lsh ; zheng2020pm , and graph-based methods malkovEfficientRobustApproximate2020 ; DBLP:journals/pvldb/WangXY021 . These ANNS indexes typically rely on heuristic algorithms to find a subset of candidates. Specifically, Inverted File Index (IVF) jegouProductQuantizationNearest2011 partitions data into clusters using the k-means algorithm and routes queries to a subset of close clusters based on the query’s distance to the clusters’ centroids. Graph-based algorithms like Hierarchical Navigable Small World (HNSW) index malkovEfficientRobustApproximate2020 construct proximity graphs and perform a beam search on the graph for a given query to find similar embeddings. Hashing-based methods like Locality-Sensitive Hashing (LSH) index DBLP:journals/corr/WangSSJ14 generate top-k candidates by hashing vectors into buckets and then retrieving the closest items from those buckets as candidates. The other type aims to accelerate the search process itself, such as quantization-based methods jegouProductQuantizationNearest2011 ; DBLP:journals/pami/GeHK014 ; gao2019beyond ; guoAcceleratingLargeScaleInference2020 ; DBLP:journals/pami/GeHK014 .
Our method belongs to the first category and is orthogonal to the second category. Our proposed index differs from previous ANNS indexes in two aspects: (1) Our proposed index is designed for embeddings based spatial keyword queries. In contrast, existing ANNS indexes are designed for embedding retrieval and do not work well when being adapted to embedding based spatial keyword queries (to be shown in Section 5.2). (2) Existing ANNS techniques are based on heuristic methods. If we want to incorporate spatial factor into the existing ANNS techniques, it is unavoidable to balance the weight between spatial and embedding relevance, which is however challenging. To avoid this, our proposed index employs neural networks to learn to cluster relevant geo-textual objects together while separating the irrelevant ones, and routes embedding based spatial keyword queries to relevant clusters through the neural networks to accelerate the top-k search process. Notably, our proposed index utilizes a novel learning-to-cluster technique to group embedding based geo-textual objects, while the IVF index employs the conventional $k$-means algorithm to cluster embeddings.
3 Problem Statement and Motivations
3.1 Preliminary
We consider a geo-textual object dataset $D$, where each geo-textual object $o\in D$ has a location description $o.loc$ and a textual description $o.doc$. The location description $o.loc$ is a two-dimensional GPS coordinate composed of latitude and longitude. The textual description $o.doc$ is a document that describes the object. The Top-k KNN Spatial-Keyword Query (TkQcongEfficientRetrievalTopk2009 is defined as follows.
Top-k KNN Spatial-Keyword Query (TkQ): Given a query $q=\langle loc,doc,k\rangle$, where $q.doc$ denotes the query keywords, $q.loc$ is the query location, and $q.k$ is the number of returned objects, we aim to retrieve $k$ objects with the highest relevance scores $ST(q,o)$:
$ST(q,o)=(1-\alpha)\times SRel(q.loc,o.loc)+$ (1)
$\alpha\times TRel(q.doc,o.doc).$
[…]
3.2 Data Analysis and Motivations
(a) Word Mismatch Analysis
(b) Spatial Relevance Analysis
Figure 1: Figure 1(a) shows the percentage distribution of ground-truth positive query-object pairs based on the number of matching terms on the Beijing dataset. Figure 1(b) compares the CDF of spatial distance for ground-truth positive query-object pairs and the linear distribution on the Beijing dataset.
Word Mismatch. As discussed in Section 1, TkQ congEfficientRetrievalTopk2009 ; DBLP:conf/icde/FelipeHR08 ; DBLP:conf/ssd/RochaGJN11 typically uses traditional information retrieval methods such as BM25 and TF-IDF DBLP:journals/ftir/RobertsonZ09 ; DBLP:journals/tois/WuLWK08 to compute textual relevance. These models rely on exact word matching to compute textual relevance and thus suffer from the word mismatch issue, which reduces their effectiveness. We further illustrate this issue in the real-world dataset that contains query-object ground-truth pairs. The details on the data are given in Section 1. As shown in Figure 1(a), in the Beijing dataset, nearly 20% of the ground-truth query-object positive (relevant) pairs have no overlap of words. This indicates that objects relevant to the query may receive a low or even zero textual relevance from these models. For instance, given a search ‘nearby drugstore’, if the search engine employs the BM25 to determine textual relevance, a drugstore labeled as ‘pharmacy’ would get a textual relevance score of zero, although the two terms convey the same meaning.
Spatial Relevance. TkQ congEfficientRetrievalTopk2009 ; liu2023effectiveness typically computes spatial relevance by $1-SDist(q.loc,o.loc)$. The implicit assumption behind this linear function is that the user’s geographical preference for geo-textual objects is a linear function of distance. However, this assumption was not examined previously and we found it does not hold on our real-life datasets. In Figure 1(b), we illustrate this issue by comparing the cumulative distribution function (CDF) of ground-truth positive (relevant) query-object pairs with the ‘Linear Distribution’, a linearly ascending hypothetical scenario that positive query-object pairs are uniformly distributed across the range $[0,1]$. This figure shows a sharp increase in the CDF of ground-truth positive pairs for $SDist(q.loc,o.loc)$ below 0.1, a pattern that greatly differs from the hypothetical scenario. This motivates us to design a new spatial relevance module to better fit the real-world scenario. An intuitive solution would be to employ an exponential function. However, our experiments indicate that this method is even less effective compared to the linear function (as to be shown in Section 5.5).
Efficiency. To address the word mismatch issue, several recent studies liu2023effectiveness ; DBLP:conf/aaai/ZhaoPWCYZMCYQ19 ; DBLP:conf/sigir/YuanLLZYZX20 have utilized deep learning techniques to evaluate textual relevance, thereby enhancing ranking effectiveness. However, these methods rely on word embeddings and complex neural networks to compute textual relevance, resulting in high querying latency. For instance, DrW liu2023effectiveness identifies the top-k relevant words from each object’s description for each query word based on their word embeddings, and then aggregates the scores of these word pairs using an attention mechanism to compute the textual relevance. On the Geo-Glue dataset, which comprises 2.8 million objects, DrW takes more than 7 seconds to answer a query on average in our experiment, which aligns with the results reported in liu2023effectiveness . This makes them unsuitable as a retriever for practical geo-textual object retrieval applications, although they can be used as re-rankers for a small number of objects returned by a retriever.
Motivations of LIST. As discussed in the last paragraph, incorporating deep textual relevance into spatial keyword queries has presented significant efficiency challenges. However, as discussed in Section 2, the first category of deep text relevance model, namely word interaction based deep relevance models, such as DrW liu2023effectiveness , which rely on word interaction modules and word embeddings to compute textual relevance, are not efficient for online computation and retrieval by a specifically designed index. Therefore, developing a new lightweight yet effective relevance model and a new index to accelerate the top-k search process for spatial keyword queries based on the proposed deep relevance model is in demand. For the first objective, we develop a lightweight yet effective deep relevance model for embedding based spatial keyword queries.
For the second objective, existing geo-textual indexes DBLP:journals/vldb/ChenCCJ21 developed for spatial keyword queries are designed based on traditional retrieval models such as BM25 and cannot be used to handle embedding based spatial keyword queries. Existing ANNS indexes, which are discussed in Section 2, do not consider spatial factors. Directly employing existing ANNS indexes for spatial keyword queries will result in severe degradation of effectiveness (shown in Section 5.2). Simply adapting existing ANNS indexes, such as the IVF index, to incorporate spatial factors will cause even more severe effectiveness degradation, as it is hard to manually set a weight to balance spatial and textual similarities (as to be shown in Section 5.2). It is still an open problem to design a new ANNS index to support embedding based spatial keyword queries.
3.3 Problem Statement
In this study, we aim to (1) design a lightweight yet effective relevance model, and (2) develop an index to expedite the top-k search process of the proposed relevance model.
4 Proposed Retriever (LIST)
Figure 2: The three phases of our retriever LIST: the training, indexing, and query phase. LIST consists of a relevance model (shown in yellow) and an index (shown in green).
4.1 Overview
LIST is featured with a new relevance model and a novel machine learning based ANNS index. Notably, our proposed index is also applicable to any other textual relevance model that transforms query keywords and object text descriptions into two separate embeddings.
The workflow of LIST is illustrated in Figure 2, which has three phases: the training, indexing, and query phase. During the training phase, we first train our relevance model, and then train our index. During the indexing phase, each object is partitioned into one cluster. During the query phase, the trained index routes each query to either a single cluster or a subset of clusters that have the highest probabilities. Within these clusters, LIST returns $k$ objects with the highest scores ranked by the trained relevance model as the query result.
4.2 The Proposed Relevance Model
Textual Relevance Learning. Inspired by the dual-encoder model’s success in document retrieval DBLP:conf/emnlp/KarpukhinOMLWEC20 , we employ a dual-encoder module to encode query keywords and object description into two seperate embeddings, and calculate textual relevance by the inner product between the two embeddings. Compared with previous relevance models liu2023effectiveness that rely on word embeddings and complex word interaction functions discussed in Section 2, the dual encoder module is both efficient and effective.
The dual-encoder module comprises an object encoder $E_{o}$ and a query encoder $E_{q}$, which are pre-trained language models like Bert devlin2018bert . The encoder takes the textual content of the query or object as input, captures interactions between words by a transformer-based model, and finally utilizes the representation of the [CLS] token as the global semantic representation, which is a $d$-dimensional embedding. This process is formulated as:
$o.emb=E_{o}(o.doc;\theta_{o}),\>o.emb\in\mathbb{R}^{d},$ (2)
$q.emb=E_{q}(q.doc;\theta_{q}),\>q.emb\in\mathbb{R}^{d}.$
where $E_{o}(\textperiodcentered;\theta_{o})$ denotes the object encoder parameterized with $\theta_{o}$ and $E_{q}(\textperiodcentered;\theta_{q})$ denotes the query encoder parameterized with $\theta_{q}$. Then the text relevance score is calculated by the inner product between the $q.emb$ and $o.emb$,
$TRel(q.doc,o.doc)=q.emb\cdot o.emb.$ (3)
Spatial Relevance Learning. As discussed in Section 3.2, it is essential to develop a more effective spatial relevance module. Therefore, we propose a new learning-based spatial relevance module to learn to estimate spatial relevance.
As discovered in Section 3.2, the user’s geographic preferences for geo-textual objects do not follow a linear pattern. Here, we consider that the user’s geographic preference for geo-textual objects increases as the distance decreases, i.e., users tend to visit nearby objects, and exhibits a stepwise decline as the distance increases. To explain this stepwise pattern, let us consider a scenario in which a customer wishes to purchase coffee from Starbucks. If the nearest Starbucks is very close, s/he would go and buy coffee. If the nearest Starbucks is a little far, s/he may hesitate and her/his intention of visiting the Starbucks would decrease. If the nearest Starbucks is very far away, s/he would give up this idea. Therefore, the new spatial relevance module is designed to be a learnable monotonically increasing step function, to align with this real-world characteristic.
Our proposed spatial relevance module takes $S_{in}=1-SDist(q.loc,o.loc)$ as input, consisting of a threshold array $\mathcal{T}\in\mathbb{R}^{t\times 1}$ and a learnable weight array $w_{s}\in\mathbb{R}^{1\times t}$. Here, $\mathcal{T}$ stores the threshold values that determine the transition points of the step function, i.e., the value exceeded by $S_{in}$ will trigger an increase of spatial relevance, which is used to ensure that the learned function is a step function. Specifically, $\mathcal{T}$ is structured as $\mathcal{T}[i]=\frac{i}{t}$, where $i\in[0,t]$ and $t$ is a hyperparameter to control the increment of the threshold value. For example, when $t=100$, $\mathcal{T}$ is $[0.0,0.01,\cdots,0.99,1.0]$. The learnable weight array $w_{s}$ determines the extent of the increase when the input $S_{in}$ reaches these threshold values, which are learned from the training data and used to estimate the spatial relevance. When the input $S_{in}$ exceeds the value of $\mathcal{T}[i]$, then the spatial relevance increases by $act(w_{s}[i])$, where $act$ is an activation function to ensure $act(w_{s}[i])$ remain non-negative. This process ensures that the output $SRel$ exhibits a step increase as the input $S_{in}$ increases. The learned spatial relevance is computed as below:
$SRel(q.loc,o.loc)=act(w_{s})\cdot\mathbb{I}(S_{in}\geq\mathcal{T}[i]),$ (4)
where $SRel(q.loc,o.loc)$ is the learned spatial relevance. $\mathbb{I}\in\{0,1\}^{t\times 1}$ is an indicator array. $\mathbb{I}[i]=1$ if $S_{in}\geq$ $\mathcal{T}[i]$; otherwise 0. The sum of the step increment is conducted by an inner product between the indicator array $\mathbb{I}\in\{0,1\}^{t\times 1}$ and the learnable weight after activation $act(w_{s})\in\mathbb{R}^{1\times t}$.
During the query phase, we extract the weights in $w_{s}$ from the module and store them as an array $\hat{w_{s}}$ for faster inference. $\hat{w_{s}}$ is constructed as $\hat{w_{s}}[i]=\sum\limits_{0}^{i}act(w_{s}[i])$. When computing the spatial relevance, we get the input $S_{in}$. Since the threshold value grows uniformly by $\frac{1}{t}$, we can determine the number of threshold values exceeded by the input as $\lfloor S_{in}*t\rfloor$, where $\lfloor$.$\rfloor$ . ⌋ indicates a floor function and is utilized to truncate a real number to an integer. This also corresponds to the sum of the values of weights, which is the spatial relevance score. This process is formulated as:
$SRel(q.loc,o.loc)=\hat{w_{s}}[\lfloor S_{in}*t\rfloor].$ (5)
Hence, during the query phase, the time complexity of computing spatial relevance is O(1), which is efficient.
Adaptive Weight Learning. The recent study liu2023effectiveness has shown the importance of weight learning in improving ranking effectiveness. To enhance ranking effectiveness, we propose an adaptive weight learning module, which aims to assign adaptive weights to textual and spatial relevance based on the query keywords. For instance, in the case of detailed query keywords, such as ‘White House, 1600 Pennsylvania Avenue NW, Washington’, giving greater weight to textual relevance better fits the real-world scenario. Hence, we employ a simple yet effective manner that directly utilizes an MLP layer to determine the weights based on the embedding of query keywords $q.emb$, which is formulated as follows:
$w_{st}=MLP(q.emb),\>w_{st}\in\mathbb{R}^{1\times 2}.$ (6)
Similar to the Equation 1, the final relevance score between the query and the object is calculated as:
$ST(q,o)=w_{st}\cdot[TRel(q.doc,o.doc),SRel(q.loc,o.loc)]^{T}.$ (7)
Training Strategy. To train our relevance model, we employ the contrastive learning strategy. Given a query $q_{i}$, the positive (relevant) geo-textual object $o_{i}^{+}$ is obtained by real-world ground-truth data (detailed in Section 5.1). As for the negative (irrelevant) geo-textual objects, following previous study liu2023effectiveness , in each training epoch, we randomly pick $b$ hard negative samples from the dataset. We optimize the loss function as the negative log-likelihood of the positive object:
$\begin{split}L_{\text{model}}&(q_{i},o_{i}^{+},o_{i,1}^{-},o_{i,2}^{-},\cdots,o_{i,b}^{-})\\ &=-log\frac{e^{ST(q_{i},o_{i}^{+})}}{e^{ST(q_{i},o_{i}^{+})}+\sum_{j=1}^{b}e^{ST(q_{i},o_{i,j}^{-})}}.\end{split}$ (8)
In addition, we also utilize the in-batch negatives strategy lin2021inbatch to further enhance training efficiency.
Figure 3: The illustration of the index.
4.3 The Proposed Index
In this section, our objective is to develop a new ANNS index that, for each embedding based spatial keyword query, can return a subset of objects that receive high scores from the proposed relevance model for the input query, thereby accelerating the top-k search process.
To build such an index, our proposed index adopts a machine learning based method. Without manually setting the weight between spatial and textual relevance, it learns to cluster relevant queries and objects together while separating the irrelevant ones. Specifically, as illustrated in Figure 3, it takes the geo-location and textual embeddings of objects or queries as input and employs a Multi-Layer Perceptron (MLP) to partition objects and route queries into $c$ clusters, and then searches for the top-k relevant objects within the routed cluster. To make the index both effective and efficient, we develop a novel pseudo-negative query-object pair generation method. These pseudo labels combined with the positive (relevant) query-object pairs ensure that relevant queries and objects are partitioned and routed into the same cluster while separating the irrelevant ones, thereby reducing the search space while maintaining effectiveness.
Feature Construction. We utilize geo-location and textual embedding to construct a consistent input representation for both objects and queries, which is illustrated in Figure 3. The textual embedding $emb\in\mathbb{R}^{d}$ is a $d$-dimensional embedding, which is converted from textual content $o.doc$ or $q.doc$ by the trained dual encoder module. The geo-location $loc=\langle lat,lon\rangle$ are transformed into the following features $\langle\hat{lat},\hat{lon}\rangle$ as below:
$\hat{lat}=\frac{lat-lat_{min}}{lat_{max}-lat_{min}},\hat{lon}=\frac{lon-lon_{min}}{lon_{max}-lon_{min}},$ (9)
where $lat_{min}$ ($lon_{min}$) represents the lowest latitude (longitude) in dataset $D$, $lat_{max}$ ($lon_{min}$) denotes the highest latitude (longitude) in dataset $D$. Then the input representation is formulated as
$x=[emb,\hat{lat},\hat{lon}].$ (10)
Here $x$ is the input feature vector for the cluster classifier, i.e., the MLP.
Cluster Classifier. Our cluster classifier is a lightweight Multi-Perception Layer (MLP). Given $x$ as the input, it produces the $c$-cluster probability distribution, defined as follows:
$Prob=Softmax(MLP(x)),\>Prob\in\left[0,1\right]^{c\times 1},$ (11)
where $c$ is a hyperparameter that indicates the desired number of clusters we aim to obtain, and $Prob$ represents the predicted $c$-cluster probability distribution. The $c$-cluster probability distribution of object $o$ is denoted as $Prob_{o}$ and that of query $q$ is represented as $Prob_{q}$.
Note that the cluster classifier, i.e., the MLP, is shared between queries and objects. This setting allows our index to learn the distribution of both queries and objects, subsequently grouping identified relevant queries and objects into the same cluster (as detailed later).
Pseudo-Label Generation. To train the cluster classifier, both positive and negative pairwise labels are required. We utilize the ground-truth query-object relevant label (e.g., click-through or human annotation, detailed in Section 5.1) as the positive pairwise label for training. However, we lack high-quality negative pairwise labels. Randomly selecting negative objects will lead to overfitting and all objects are grouped together (as detailed later). Therefore, we propose a novel pseudo pairwise negative label generation method.
The relationship between query $q$ and object $o$ is denoted as $s(q,o)$. If $q$ and $o$ are relevant , $s(q,o)=1$; otherwise, $s(q,o)=0$, which are then used as labels for training. We leverage ground-truth positive labels as pairwise positive labels in the training, as shown below:
$\text{pos}_{q}=\{o;o\in D,s(q,o)=1\},$ (12)
where the positive object set of query $q$ is denoted as $\text{pos}_{q}$.
The negative object set $\text{neg}_{q}$ of query $q$ is generated by the relevance model. Given a query $q$, we employ the relevance model to calculate the relevance score for all objects. Then we adaptively select $\text{neg}_{q}$ according to two hyperparameters $neg_{start}$ and $neg_{end}$, as shown below:
$\text{neg}_{q}=\text{argsort}_{o\in D}ST(q,o)[neg_{start}:neg_{end}],\;s(q,o)=0,$ (13)
where $ST(q,o)$ is the relevance score between query $q$ and object $o$ produced by the relevance model. Note that positive query-object pairs are excluded as indicated by the filter condition $s(q,o)=0$.
This adaptive pseudo-negative generation method draws inspiration from hard negative sample strategy DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 . What sets our approach apart from existing studies DBLP:conf/naacl/QuDLLRZDWW21 ; renRocketQAv2JointTraining2021 is that we employ an adaptive manner to select hard negative samples, thereby controlling the difficulty level of the generated negative samples. This adjustment strikes the trade-off between the effectiveness and efficiency.
Decreasing $neg_{start}$ leads to a set of harder negative objects being chosen to train the model. Consequently, the classifier is more effective in distinguishing between positive objects and hard negative objects, which facilitates the clustering of relevant queries and objects, while effectively segregating the irrelevant ones. Empirically, under this setting, only the very relevant queries and objects are grouped into the same cluster. The reduced number of objects within a query’s cluster leads to higher efficiency. In contrast, when using a large $neg_{start}$, the index can not learn useful information, leading to a scenario where all objects tend to be clustered together.
On the other hand, when our index is trained on a set of harder negative objects, it might also exclude some positive ones and reduce its effectiveness. Thus, the choice of $neg_{start}$ strikes a trade-off between effectiveness and efficiency. This alleviates the lack of negative label issue and the skewed cluster distribution issue of existing techniques Hsu16_KCL ; Hsu18_L2C .
Training Strategy. Based on the hard negative objects and positive objects provided above, we employ the MCL loss function Hsu19_MCL to train the MLP, as described below:
$\begin{split}L_{\text{Index}}&(q_{i},o_{i}^{+},o_{i,1}^{-},o_{i,2}^{-},\cdots,o_{i,m}^{-})\\ &=log(\hat{s}(q_{i},o_{i}^{+}))+\sum\limits_{j=1}^{m}log(1-\hat{s}(q_{i},o_{i,j}^{-})),\end{split}$ (14)
where $\hat{s}(q_{i},o_{j})=Prob^{T}_{q_{i}}\cdot Prob_{o_{j}}$, $o_{i}^{+}\in\text{pos}_{q}$ and $o_{i,j}^{-}\in\text{neg}_{q}$. Typically, we randomly select one positive object $o_{i}^{+}$ from the positive object set $\text{pos}_{q}$ and $m$ negative objects from the negative object set $\text{neg}_{q}$ in each training epoch.
As described earlier, the MLP is shared between queries and objects. Through this training process, for positive query-object pairs, their $Prob_{o}$ and $Prob_{q}$ will have a similar distribution. Thereby, relevant pairs of query and object are more likely to be grouped into the same cluster while the irrelevant pairs are grouped into distinct clusters. For example, as illustrated in Figure 3, the object $o_{4}$ and query $q_{2}$ form a positive query-object pair, and they are expected to be grouped into the same cluster $C_{2}$.
Learning to Partition and Route. During the indexing phase, each object $o$ is partitioned to the cluster with the highest probability according to $Prob_{o}$. Once the partitioning is completed, the objects assigned to a cluster are stored in a corresponding list, which acts as an inverted file for these objects. Each object in this list is represented by a $d$-dimensional vector $o.emb$, and a geo-location $o.loc$, which will be utilized to calculate the relevance score for incoming queries. During the query phase, a given query $q$ is directed to the cluster that has the highest probability. Subsequently, the relevance scores between $q$ and all objects within the cluster are calculated and the top-k objects are selected as result objects.
An alternative way is routing queries (objects) to $cr$ clusters with the highest probabilities based on $Prob_{q}$ ($Prob_{o}$). Although this might boost query effectiveness by considering more objects in different clusters, it sacrifices efficiency as more objects need to be calculated for a query, leading to an accuracy-efficiency trade-off.
Insertion and Deletion Policy. When a new object comes, we convert it into an embedding using the trained relevance model and then assign it to specific clusters using the trained index. When an object is deleted, we simply remove it from the corresponding cluster. Note that most existing ANNS indexes are static luoSurveyDeepHashing2023 ; DBLP:journals/pami/WangZSSS18 ; DBLP:journals/pvldb/WangXY021 . When inserted data significantly differs in distribution from the existing data or when insertions occur frequently, periodically rebuilding the index is necessary to maintain high accuracy. Our index follows this approach. Apparently, the time costs of insertion and deletion operations are negligible.
Cluster Evaluation. After the training phase, we employ the trained index to produce clusters $C=\{C_{1},C_{2},\\ \cdots,C_{c}\}$, where $C_{i}$ represents cluster $i$. We use validation queries to evaluate the quality of the clusters. The validation queries are fed into the trained index and routed to a cluster. $C_{i}^{q}$ denotes a list containing the validation queries routed to $C_{i}$, $C_{i}^{o}$ denotes a list containing objects partitioned to $C_{i}$, and ||||| | indicates the size of a list. We proceed to introduce two metrics to evaluate the quality of the clusters.
The first metric evaluates the precision of $C_{i}$, denoted as $P(C_{i})$, representing the degree to which queries are aligned with their corresponding positive (relevant) objects in the same cluster, which is defined below:
$P(C_{i})=\frac{1}{|C_{i}^{q}|}\sum\limits_{q_{j}\in C_{i}^{q}}\frac{|\text{pos}_{q_{j}}\cap C_{i}^{o}|}{|\text{pos}_{q_{j}}|}.$ (15)
Building upon this, we compute the average precision across all clusters, denoted as $P(C)$, which is defined as below:
$P(C)=\frac{1}{\sum\limits_{C_{i}\in C}|C_{i}^{q}|}\sum_{C_{i}\in C}P(C_{i})*|C_{i}^{q}|.$ (16)
Intuitively, a higher $P(C)$ indicates that the index is more effective.
In addition, we also take into account efficiency concerns. For this purpose, we introduce another metric, the Imbalance Factor (IF) johnson2019billion , which measures the degree of balance across all clusters, denoted as $\text{IF}(C)$. The $IF(C)$ is formulated as:
$\text{IF}(C)=\frac{\sum|C_{i}|^{2}}{(\sum|C_{i}|)^{2}},$ (17)
where $\text{IF}(C)$ is minimized when $|C_{1}|=|C_{2}|=\cdots=|C_{c}|$ according to the Cauchy-Schwarz Inequality. A higher imbalance factor indicates a more uneven distribution of clusters. When most objects are concentrated in a few large clusters, the imbalance factor increases significantly, which is undesirable for our task. Overall, our goal is to achieve high-quality clustering results characterized by higher $P(C)$ and lower $\text{IF}(C)$. In our index, with proper hyperparameters, our training process of the cluster classifier is able to obtain high-quality clusters. In the experiments, We study the quality of the clusters produced by our proposed index using these evaluation metrics (shown in Section 5.5).
4.4 Procedures and Analyses of LIST
Input: A geo-textual object dataset $D$, a relevance model $R$, a index $I$, training TkQs set $Q_{\text{train}}$, and incoming TkQs set $Q$
Output: The response $Res_{q}$ for each query $q\in Q$
1 // Training Phase: $Train(Q_{\text{train}},D,R,I)$
2 Train $R$ by $Q_{\text{train}}$, $D$ based on Eqaution 8;
3 Employ $R$ to generate pseudo-labels based on Equation 13;
4 Employ pseudo-labels to train $I$ based on Equation 14;
5 // Indexing Phase: $Indexing(D)$
6 for $o\in D$ do
7       Transform $o.doc$ to $o.emb$ based on Equation 2;
8       Transform $o.emb$, $o.loc$ to $x_{o}$ based on Equation 10;
9       Generate $Prob_{o}$ by $MLP$ based on Equation 11;
10       Parition $o$ to cluster $C_{i}$ based on $Prob_{o}$;
11
12 // Query Phase: $Search(q,C,R,I)$
13 for $q\in Q$ do
14       Transform $q.doc$ to $q.emb$ based on Equation 2;
15       Transform $q.emb$, $q.loc$ to $x_{q}$ based on Equation 10;
16       Generate $Prob_{q}$ by $I$ based on Equation 11;
17       Route $q$ to cluster $C_{i}$ based on $Prob_{q}$;
18       $Res_{q}\leftarrow\text{argTop-k}_{o\in C_{i}}ST(q,o)$;
19       return $Res_{q}$;
20
Algorithm 1 Procedures of LIST.
Procedures of LIST. The detailed procedures of LIST are summarized at Algorithm 1. In the training phase, we train the proposed relevance model and index (lines 2-4). After that, in the indexing phase, we assign all objects to their corresponding clusters as inverted files (lines 6-10). In the query phase, given a new query $q$, we extract its features and then route it to a cluster $C_{i}$ by the index (lines 13-16). Consequently, we calculate the relevance score between $q$ and each object $o$ in that cluster by the relevance model, and the top-$k$ objects with the highest scores are returned to answer query $q$ (lines 17-18).
Complexity Analysis. Now we analyze the time and space complexity of LIST. Assuming the dimensionality of the embedding is $d$, the total number of objects is $n$, and the number of layers in the MLP of the index is $l$. The embeddings $o.emb$ of objects are generated in advance.
For a given query, the time complexity of LIST is $O((l-2)d^{2}+dc+\frac{n}{c}(d+2)+d)$. $O((l-2)d^{2}+dc)$ is the time complexity for our index (lines 15), which is the inference computation cost of Equation 11. $O(\frac{n}{c}(d+2)+d)$ is the time cost of our relevance model (line 17). $\frac{n}{c}$ denotes the number of objects to be checked by the relevance model, which is approximately $\frac{1}{c}$ of the entire dataset. This is because of the even cluster distribution (Verified in Section 5.5), which means approximately $\frac{1}{c}$ of the dataset needs to be visited. Specifically, $O(\frac{nd}{c})$ is the computational cost of Equation 3. $O(\frac{2n}{c})$ is the computational cost of Equation 5, and $O(d)$ is the computational cost of Equation 6.
The space complexity of LIST is $O((l-1)d+dc)+n(d+2)$. $O((l-1)d+dc)$ is the size of the MLP used by our index. $O(nd)$ represents the space required for storing pre-computed $o.emb$, and $O(2n)$ denotes the storage cost for geo-locations.
5 Experiments
Table 1: Datasets Statistics.
Beijing
Shanghai
Geo-Glue
Number of Pois
122,420
116,859
2,849,754
Number of Queries
168,998
127,183
90,000
Number of Records
233,343
182,634
90,000
Number of Train Queries
136,890
103,019
50,000
Number of Val Queries
15,209
11,446
20,000
Number of Test Queries
16,899
12,718
20,000
Number of Train Records
189,027
148,017
50,000
Number of Val Records
21,034
16,492
20,000
Number of Test Records
23,282
18,125
20,000
We evaluate the effectiveness and efficiency of our proposed solution in answering TkQs by comparing with state-of-the-art methods on three real-world datasets. We aim to answer the following research questions:
  • RQ1: Does our proposed retriever LIST and our proposed relevance model outperform previous state-of-the-art methods in terms of effectiveness?
  • RQ2: How well does the efficiency of our proposed retriever LIST compare with state-of-the-art methods?
  • RQ3: How does LIST scale with the size of dataset?
  • RQ4: What are the impacts of our proposed index and different modules in the proposed LIST?
5.1 Experimental Setup
Datasets. To evaluate the effectiveness and efficiency of our proposed retriever LIST, we utilize three benchmark datasets: Beijing, Shanghai, and Geo-Glue. The Beijing and Shanghai datasets liu2023effectiveness are provided by Meituan, a Chinese retail services platform. Users submit a query through the platform, which consists of a query location and a set of keywords. Subsequently, the Points of Interest (POIs) that users clicked on are recorded in the query log, and is considered as a ground-truth. Note that the click-through data recorded in the search log may be the only feasible way to get a large scale of ground truth data for spatial keyword queries liu2023effectiveness . The explicit feedback such as ratings is very challenging to collect DBLP:conf/www/HeLZNHC17 . Note that using query logs is also the popular way of generating ground truth in the Information Retrieval literature DBLP:conf/www/HeLZNHC17 ; DBLP:conf/cikm/HuangHGDAH13 ; DBLP:conf/www/YaoTCYXD021 ; DBLP:conf/kdd/Joachims02 ; DBLP:journals/csur/WuSZXC23 . Therefore, following the previous work liu2023effectiveness , we utilize the two datasets to evaluate our proposed solution LIST and treat the clicked POIs as ground truth relevant objects to the corresponding queries. The Geo-Glue dataset li2023geoglue ; DBLP:conf/sigir/DingCXHLZX23 is released by DAMO Academic. The POIs in the Geo-Glue dataset are crawled from OpenStreetMap [^1] https://www.openstreetmap.org/, and the queries and the corresponding ground truth POIs are manually generated by domain experts. Notably, in the released Geo-Glue dataset, the coordinates of objects and queries are modified due to privacy considerations, which results in many objects with identical geo-locations.
The release of query log datasets from the industry is highly restricted. As a result, to the best of our knowledge, there are no other public datasets that contain ground-truth query results or query logs, except for the three datasets used in our experiments. To investigate the scalability of the proposed framework, we conduct a scalability study to show LIST’s efficiency on larger datasets, where the Geo-Glue dataset is augmented with more crawled POIs (shown in Section 5.4)
Dataset Split. The statistics of datasets are stated in Table 1, where each record represents a single ground-truth label between an object and a query, and each query may have multiple ground-truth objects. For the Beijing and Shanghai datasets, to ensure a fair comparison, we follow the previous split strategy liu2023effectiveness , where 90% of queries and their corresponding ground-truth records are used as the training set and the remaining queries as the test set. From the training set, we randomly choose 10% data as the validation set to tune hyperparameters. For the Geo-Glue dataset DBLP:conf/sigir/DingCXHLZX23 , we follow the provided splits for training, validation, and test data.
Table 2: Comparison of relevance models across three datasets by brute-force search.
Beijing Shanghai Geo-Glue
Recall NDCG Recall NDCG Recall NDCG
@20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
TkQ
0.5740
0.5283
0.4111
0.3302
0.6746
0.6380
0.5044
0.4009
0.5423
0.5023
0.3847
0.3051
PALM
0.3514
0.3098
0.2077
0.1343
0.4617
0.4023
0.2065
0.1223
N/A
N/A
N/A
N/A
DrW
0.6968
0.6316
0.4814
0.3791
0.7689
0.7159
0.5394
0.4114
N/A
N/A
N/A
N/A
DPR
0.5078
0.4269
0.2892
0.2209
0.5143
0.4161
0.2505
0.1721
0.7626
0.7249
0.6014
0.5079
${}_{D}$
0.7555 0.6845 0.5181 0.4215 0.7986 0.7494 0.5769 0.4568 0.7963 0.7565 0.6211
0.5202
MGeo
N/A
N/A
N/A
N/A
N/A
N/A
N/A
N/A
0.7049
N/A
N/A
0.5270
LIST-R
0.8156 0.7545 0.5913 0.4989 0.8361 0.7924 0.6445 0.5397 0.8393 0.8033 0.6837 0.5887
7.95%
10.22%
14.12%
18.36%
4.69%
5.73%
11.71%
18.14%
5.73%
6.18%
6.26%
11.70%
Effectiveness metric. Following previous studies DBLP:conf/sigir/DingCXHLZX23 ; liu2023effectiveness , we use two metrics, i.e., the Recall and Normalized Discounted Cumulative Gain (NDCG), to evaluate the effectiveness. Recall@$k$ evaluates the proportion of positive objects contained in the top-$k$ candidates for a given query. NDCG@$k$ considers the order of ground-truth objects in the retrieved objects, reflecting the quality of the ranking in the retrieved list. We assign the graded relevance of the result at position $i$ as $rel_{i}\in\{0,1\}$, where $rel_{i}=1$ when the object is relevant to the query, otherwise $rel_{i}=0$. More details can be found in liu2023effectiveness . Specifically, the following metrics are utilized: Recall@10, Recall@20, NDCG@1, and NDCG@5.
Baselines of Relevance Model. We compare our proposed relevance model, denoted as LIST-R, with various state-of-the-art spatio-textual relevance models or state-of-the-art deep textual relevance models. We also modify the state-of-the-art textual relevance model DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 into ${}_{D}$ to answer TkQs.
[…]
  • PALM DBLP:conf/aaai/ZhaoPWCYZMCYQ19 : This method employs deep neural networks for query-object spatio-textual relevance.
  • DrW liu2023effectiveness : This is the newest deep relevance based method for answering TkQs.
  • DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 : This is a state-of-the-art textual relevance model for document retrieval. This method does not consider spatial relevance.
  • ${}_{D}$: This method replaces the text relevance computation in DrW liu2023effectiveness with DPR DBLP:conf/emnlp/KarpukhinOMLWEC20 .
  • MGeo DBLP:conf/sigir/DingCXHLZX23 : This is a recent deep relevance based method for answering TkQs. Note that we cannot reproduce the experimental results reported in the paper by running the official code, and we use the evaluation results from the original paper.
[…]
An important hyperparameter is $k$, which represents the number of objects retrieved by the indexes and reranks by ${}_{D}$. Specifically, on the Beijing and Shanghai datasets, for a fair comparison, we set $k$ to be 10,000, which is the average number of objects retrieved by our index per query. On the geo-glue dataset, $k$ is set to be 55,000 for the same reason above. We evaluate the effect of $k$ in Section 5.3.
  • TkQ+${}_{D}$: This method employs TkQ to retrieve top-$k$ objects and then reranks these objects by ${}_{D}$.
  • IVF+${}_{D}$: This method constructs an IVF index over the embeddings produced by ${}_{D}$, and the objects within the selected index cluster are reranked by ${}_{D}$. Notably, this method requires two parameters: the number of clusters $c$, and the number of clusters to route for each query and object $cr$. It does not involve $k$. We set $c$ and $cr$ to be the same as our method across three datasets for a fair comparison.
[…]
  • LSH+${}_{D}$: This method constructs an LSH index with the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by fetching similar embeddings in the same buckets, and reranks the retrieved objects by ${}_{D}$. Following previous work johnson2019billion , we set the length of the hash code $nbits$ to 128.
  • HNSW+${}_{D}$: This method constructs an HNSW index over the embeddings produced by ${}_{D}$, which retrieves the top-$k$ relevant objects by conducting beam searches over the proximity graph, and reranks the retrieved objects by ${}_{D}$. Following previous work malkovEfficientRobustApproximate2020 , we set the number of links $M$ to 48 and $efConstruction$ to 100.
  • IVFPQ+${}_{D}$: This method integrates the IVF index with the product quantization technique jegouProductQuantizationNearest2011 to retrieve objects. The retrieved objects are reranked by ${}_{D}$. We set the number of clusters $c$ to be the same as our index. Following the instruction johnson2019billion , we set the number of centroids $w$ to 32 and the number of bits $nbits$ to 8, the number of clusters to search $cr$ at 4. Here, we set a larger value for $cr$ than that in our index since the product quantization accelerates computations within clusters, allowing access to more clusters with comparable time costs. Notably, the maximal $k$ supported by the Faiss library is 2,048, therefore we set $k$ to 2,048 across three datasets.
Table 3: Comparison of retrievers across three datasets.
Beijing Shanghai Geo-Glue
Recall NDCG Recall NDCG Recall NDCG
@20 @10 @5 @1 @20 @10 @5 @1 @20 @10 @5 @1
TkQ+${}_{D}$
0.7263 0.6633 0.5056 0.4145 0.7717 0.7348 0.5742 0.4556 0.7357 0.7012 0.5797 0.4878
IVF+${}_{D}$
0.5461
0.4954
0.3771
0.3103
0.4965
0.4717
0.3713
0.3000
0.4720
0.4535
0.3835
0.3274
[…]
0.4126
0.3714
0.3025
0.2415
0.4174
0.3812
0.2960
0.2101
0.3921
0.3714
0.3011
0.2419
LSH+${}_{D}$
0.5473
0.5057
0.3973
0.3385
0.5356
0.5097
0.4135
0.3453
0.6686
0.6433
0.5445
0.4661
HNSW+${}_{D}$
0.6267
0.5828
0.4545
0.3765
0.5527
0.5270
0.4182
0.3432
0.5009
0.4807
0.4066
0.3468
IVFPQ+${}_{D}$
0.7064
0.6400
0.4872
0.4002
0.7184
0.6765
0.5257
0.4188
0.6997
0.6724
0.5650
0.4791
LIST 0.7711 0.7170 0.5668 0.4812 0.7721 0.7401 0.6099 0.5139 0.8250 0.7909 0.6747 0.5815
5.63%
7.95%
10.70%
15.32%
0.05%
0.72%
6.21%
12.54%
12.13%
12.79%
16.38%
19.21%
Implementations. The relevance model and the index are trained using Pytorch. During the query phase, the index and all relevance models are inferred in C++ by the ONNX system [^3] https://github.com/onnx/onnx. In our model, we utilize the bert-base-Chinese pre-trained model [^4] https://huggingface.co/bert-base-chinese from the huggingface Library wolf2019huggingface as encoders, which is the same as previous work liu2023effectiveness ; DBLP:conf/sigir/DingCXHLZX23 . The footstep controller $t$ of $\mathcal{T}$ is set to be 1000 for all three datasets. $\text{neg}_{\text{start}}$ is set to 50,000 for the Beijing dataset, 60,000 for the Shanghai dataset, and 200,000 for the Geo-Glue dataset. We empirically set the cluster number $c$ to approximately $\frac{n}{10,000}$, i.e., 20 for the Beijing and Shanghai datasets and 300 for the Geo-Glue dataset. This is because ranking 10,000 objects by the relevance model is computationally feasible and does not notably compromise effectiveness. The number of clusters to route $cr$ for queries and objects is set to 1 by default, with different $cr$ and $c$ settings shown in Section 5.3. The implementations of DrW are from publicly available source codes, and we make use of the implementation of IVF, LSH, IVFPQ, and HNSW provided by the Faiss library johnson2019billion , while others are implemented by ourselves. The Faiss library is implemented in C++, providing a fair comparison with our index. Our default experiment environment is CPU: Intel(R) Core(TM) i9-10900X CPU@3.70GHz, Memory: 128G, and GPU: RTX3080 10GB.
5.2 Effectiveness Study (RQ1)
Effectiveness of Relevance Model LIST-R. To validate our proposed relevance model’s effectiveness, denoted as LIST-R, we compare it with other relevance models across the three datasets. All relevance models perform a brute-force search over the entire dataset to identify the top-k objects. Table 2 reports the effectiveness of the evaluated methods. DrW and PLAM cannot be evaluated on the Geo-Glue dataset via brute-force search because of their slow querying speed, requiring more than a day for evaluation. We have the following findings: (1) LIST-R consistently outperforms all the baseline models on all three datasets across every metric. Specifically, LIST-R achieves up to a 19.21% improvement over the best baseline on NDCG@1 and up to a 12.79% improvement on recall@10. (2) Traditional ranking methods are less effective than deep relevance models. On the three datasets, TkQ is outperformed by DrW and ${}_{D}$. This can be attributed to the word mismatch issue discussed in Section 3.2. Notably, PALM is outperformed by the TkQ, which is consistent with the results reported by liu2023effectiveness . (3) Directly applying the dual-encoder model to our task is less effective. On the Beijing and Shanghai datasets, DPR is outperformed by TkQ, which is caused by the ignorance of spatial relevance. DPR’s enhanced effectiveness on the geo-glue dataset can be attributed to the property of the dataset, where most queries are address descriptions DBLP:conf/sigir/DingCXHLZX23 .
[…]
Table 4: Query runtime on three datasets (ms).
Beijing
Shanghai
Geo-Glue
TkQ
6.3
6.5
1,178.0
TkQ+${}_{D}$
86.3
84.9
5,412.3
IVF+${}_{D}$
2.9 2.9 4.0
LSH+${}_{D}$
5.2
5.1
31.1
HNSW+${}_{D}$
7.3
6.5
14.6
IVFPQ+${}_{D}$
3.6
3.7
5.2
LIST 3.0 3.0 5.1
5.3 Efficiency Study (RQ2)
Runtime of LIST. We investigate the querying speed of LIST by comparing it with the state-of-the-art baselines. During the query phase, our proposed index, IVF, LSH, IVFPQ, and HNSW remain in the GPU, while IR-Tree remains in the memory. The total runtime is computed by calculating the average time cost of 5,000 randomly sampled queries from the test query set. We present the total runtime of different methods in Table 4. Based on the experimental results, we have the following findings: (1) Compared with the most effective baselines, i.e., TkQ+${}_{D}$, LIST is one order of magnitude faster on the Beijing and Shanghai datasets and three orders of magnitude faster on the Geo-Glue dataset. (2) IVF+${}_{D}$ and IVFPQ+${}_{D}$ have similar runtimes as LIST. However, their effectiveness is much worse than LIST as stated in Table 3. Specifically, compared with the IVFPQ+${}_{D}$, LIST achieves an improvement of up to 22.70% on the NDCG@1 metric across the three datasets. Compared with the IVF+${}_{D}$ baseline, we observe an improvement of up to 77.61% on the NDCG@1 metric across the three datasets.
(a) Beijing Dataset
(b) Beijing Dataset
(c) Shanghai Dataset
(d) Shanghai Dataset
Figure 4: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters ($c$) (up and right is better).
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters ($c$). We examine LIST’s ability to trade off effectiveness and efficiency under the different number of cluster $k$ settings. Specifically, on the Beijing and Shanghai datasets, we set the hyperparameter cluster number $c$ of LIST, IVF+${}_{D}$, and IVFPQ+${}_{D}$ from 5, 10, 20, 50, to 100. For other baselines, i.e., LSH+${}_{D}$, HNSW+${}_{D}$, and TkQ+${}_{D}$, we set hyperparameter the number of objects retrieved (top-$k$) from 40,000, 20,000, 1,$0$, 5,000 to 2,000. This ensures the number of objects retrieved for reranking is consistent across different baselines, leading to a fair comparison. Other hyperparameter settings are fixed. We present the trade-off results on the NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 4 while other metrics and datasets have similar results. We have the following findings: (1) LIST outperforms other methods on both effectiveness and efficiency. This demonstrates LIST’s capability of maintaining effectiveness while accelerating querying speed. (2) The effectiveness of TkQ+${}_{D}$ is consistent as $k$ increases while time overhead increases significantly. This is due to the word mismatch issue discussed in Section 3.2. Many relevant objects without word overlap with the query remain hard to retrieve as $k$ increases. (3) HNSW+${}_{D}$ exhibits almost constant results. This is because HNSW builds a sparse proximity graph, limiting the reachable objects for each query. Therefore, when $k$ exceeds the number of reachable objects, its performance remains unchanged.
(a) Beijing Dataset
(b) Beijing Dataset
(c) Shanghai Dataset
(d) Shanghai Dataset
Figure 5: The Effectiveness-Speed trade-off curve varies with the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$) (up and right is better).
Trade-off Study by varying the number of objects retrieved (top-$k$) and the number of clusters to route ($cr$). We further investigate the impact of the number of clusters $cr$ to route queries and objects. We vary $cr$ from 1, 2, 3, to 4 for IVF+${}_{D}$ and our index, from 4, 5, 6, to 7 for IVFPQ+${}_{D}$. For other baselines, we set hyperparameter the number of objects retrieved (top-$k$) from 10,000, 20,000, 3,$0$, to 4,$0$ for a fair comparison like the previous trade-off study. Other hyperparameter settings are fixed. Similarly, we report the trade-off results on NDCG@1 and Recall@10 metrics for the Beijing and Shanghai datasets in Figure 5. According to the experiment results, LIST outperforms other methods in both effectiveness and efficiency under different $cr$ settings. Increasing $cr$ can enhance effectiveness, but it also leads to higher latency. In practice, users can adjust this parameter according to their needs.
Memory Consumption of LIST. The memory consumption of LIST is composed of three parts: the memory used by the proposed relevance model, the memory used by the proposed index, and the memory used for object text embeddings that are produced in advance. The experiment results are presented in Table 5, which demonstrates the remarkable memory efficiency of LIST. Compared with LSH+${}_{D}$, TkQ+${}_{D}$, and HNSW+${}_{D}$, LIST requires less memory. The reason is that our index stores only a lightweight MLP $c$-cluster classifier. This storage requirement is less than the memory consumption of the proximity graph of HNSW, hash tables of LSH, and the inverted file of the IR-Tree.
Table 5: Memory usage on three datasets (MB).
Beijing
Shanghai
Geo-Glue
TkQ+${}_{D}$
719
857
12,037
IVF+${}_{D}$
508
505
8,510
LSH+${}_{D}$
513
511
8,638
HNSW+${}_{D}$
548
545
9,427
IVFPQ+${}_{D}$
508
505
8,520
LIST
508
505
8,515
5.4 Scalibility Study (RQ3)
Scalability Study. We further explore the scalability of LIST and the proposed relevance model LIST-R. We supplement the Geo-Glue dataset with new POIs crawled from Open Street Maps in Hangzhou. Since there is no ground-truth data for the relevance between the newly crawled POIs and the existing queries, we only report the efficiency of our proposed retriever LIST and our relevance model LIST-R. For the newly acquired POIs, we utilize the trained dual-encoder module and index to partition them into distinct clusters. We report the runtime of LIST and LIST-R on Figure 6. We observe that, as the number of objects increases, the runtime of LIST and LIST-R scales linearly.
(a) Scalability Study of LIST-R
(b) Scalability Study of LIST
Figure 6: Scalability study on Geo-Glue dataset.
5.5 Ablation Studies (RQ4)
Index Study. To validate the contribution of our proposed index to the overall effectiveness, we compare it with other baseline indexes, including TkQ, IVF, LSH, and HNSW on the Geo-Glue dataset. We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects. The results are reported in Table 6, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings. We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table 4), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1. (2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed. The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table 4.
Table 6: Index ablation study on the Geo-Glue dataset.
Recall NDCG
@20 @10 @5 @1
TkQ+LIST-R
0.7890
0.7585
0.6517
0.5639
IVF+LIST-R
0.5381
0.5192
0.4671
0.3960
LSH+LIST-R
0.7665
0.7381
0.6378
0.5555
HNSW+LIST-R
0.7527
0.7231
0.6219
0.5390
LIST 0.8250 0.7909 0.6747 0.5815
Cluster Quality Study. To illustrate the quality of produced clusters by our index, we conduct a comparison study. We use the proposed relevance model to generate embeddings and then employ our index (denoted as LIST-I) and IVF index to produce clusters separately. We present the cluster results of $IF(C)$ and $P(C)$ in Table 7, which shows that our index achieves much higher precision and obtains comparative imbalance factors compared with IVF index.
Table 7: Comparison of the quality of clusters.
Beijing Shanghai
IF(C)
$P$(C)
IF(C)
$P$(C)
IVF
1.31
0.6774
1.33
0.6418
LIST-I
1.49
0.8907
1.43
0.8382
Pesudo-Label Study. As discussed in Section 4.3, the parameter $\text{neg}_{\text{start}}$ affects the difficulty level of pseudo-negative labels, which then impacts the effectiveness and efficiency of our index. To investigate the impacts of the pseudo-negative labels, we vary the hyperparameter $\text{neg}_{\text{start}}$ from 40,000, 50,000, 60,000, and 70,000, to 80,000 on the Beijing and Shanghai datasets. We illustrate the metrics $P(C)$ and $\text{IF}(C)$ of produced clusters in Figure 7. Notably, as $\text{neg}_{\text{start}}$ increases, both $IF(C)$ and $P(C)$ tend to increase. An increased $IF(C)$ suggests a more concentrated distribution of objects, while an increased $P(C)$ indicates improved accuracy in the retrieval results. The experiment results indicate that the choice of $\text{neg}_{\text{start}}$ leads to a trade-off between effectiveness and efficiency, which can be set flexibly in real-world applications.
(a) The impact over $IF(C)$
(b) The impact over $P(C)$
[…]
Spatial Learning Study. To evaluate the learning-based spatial relevance module, we consider the following baseline: (1) LIST-R+$S_{in}$ that replaces the learning-based spatial relevance module with $S_{in}$ for training, where $S_{in}$ = 1 - $SDist(q.loc,o.loc)$; and (2) LIST-R+$\alpha*S_{in}^{\beta}$ which substitutes the spatial relevance module with a learnable exponential function. $\alpha$ and $\beta$ are two learnable parameters and are processed to ensure non-negative. Table 8 presents the experimental results obtained by conducting a brute-force search on the Beijing dataset using the trained models. Notably, LIST-R outperforms LIST-R+$S_{in}$ in all metrics. Interestingly, the first variant outperforms the second variant, which suggests that without careful design, a learnable function may be outperformed by a simple approach.
Table 8: Ablation study of spatial relevance module and weight learning via brute-force search on the Beijing dataset.
Recall NDCG
@20 @10 @5 @1
LIST-R + $S_{in}$
0.7526
0.7087
0.5255
0.4271
LIST-R + $\alpha*S_{in}^{\beta}$
0.5308
0.4532
0.3130
0.2411
LIST-R + ADrW
0.7925
0.7414
0.5832
0.4792
LIST-R
0.8156 0.7545 0.5913 0.4989
Weight Learning Study. We conduct a comparison experiment between our weight learning module and the attention mechanism proposed by DrW liu2023effectiveness (denoted as ADrW). Table 8 reports the results obtained by conducting a brute-force search on the Beijing dataset using the trained relevance models. Here, LIST-R represents our weight learning mechanism, while LIST-R+ADrW denotes replacing it with the ADrW for training. The results indicate that our weight learning mechanism surpasses the latest ADrW mechanism.
Effect of Training Dataset Size on Effectiveness. We investigate the impact of training dataset size over LIST and LIST-R. Specifically, we exclude a certain percentage of objects along with their corresponding records during the training and validation process. During testing, we use the complete test dataset. This approach avoids the issue of data leakage. We vary the percentage of objects from 30%, 50%, 80% to 100%. We only present the results for NDCG@1 on the Beijing dataset, as the remaining metrics and datasets have similar results. The results are shown in Figure 8(a). LIST-R uses brute-force search with the relevance model, while LIST utilizes our proposed index for retrieval. The performance gap between them is consistently small, which underscores the capability of our index to boost retrieval speed without sacrificing effectiveness.
(a) The Impact over Effectiveness
(b) The Impact over Training time
Figure 8: The impact of training dataset size over effectiveness and training time on the Beijing dataset.
In addition, even with low training percentages (e.g., 0.3), our proposed method maintains satisfactory effectiveness, demonstrating its ability to learn from limited data and adapt to new queries.
Effect of Training Dataset Size on Training Time. We also measure the training time per epoch of LIST-R and LIST on the Beijing dataset by varying the size of the training data in the same manner, and the results are shown in Figure 8(b). Other datasets exhibit similar trends. We observe that the training time of LIST-R and LIST appears to be linear to the size of the training data, aligning with our complexity analysis.
6 Conclusions and Future Work
In this paper, we propose a novel retriever LIST, which comprises a new relevance model and a novel index. We conduct extensive experiments to show the effectiveness and efficiency of LIST over the state-of-the-art baselines. This work opens up a promising research direction of designing novel ANNS indexes for accelerating the search for embedding based spatial keyword queries. One interesting future direction is integrating our index with product quantization techniques to further expedite the search process. Another potential direction is to extend our proposed index to vector databases for dense vectors without spatial information.